xx KadathThorn

KadathThorn

Samuel Tootle <tootle@itp.uni-frankfurt.de>

February 19, 2023

Abstract

This thorn provides the base interface to the Frankfurt University/KADATH (FUKA) library. For details on FUKA and the initial data import, see the KadathImporter thorn documentation.

1 Introduction Importing FUKA initial data currently requires linking the EinsteinToolkit executable to the built FUKA library. This thorn handles linking to an existing library or building one from scratch.

2 Linking

Linking FUKA to the ETK is done in two ways, linking to an existing FUKA library or building one during the ETK build process and linking after.

Existing library

To build using an existing library, one can set the KADATHxDIR variable in the ETK build config file to the base directory of the cloned FUKA distribution.

Alternatively, during the FUKA installation procedure, the base directory must be included in the HOMExKADATHxenvironment variable. In the event KADATHxDIR is not set, the build script will check HOMExKADATHx for a previously built library. If one is found, this library will be used. If one wishes to force building FUKA during the ETK build procedure, they can set

xxxxKADATH_DIRx=xBUILD

in the ETK installation configuration.

Build

If a previously built version of the FUKA library is not found, a manual installation is done during the ETK build procedure. The FUKA library is a git submodule of the KadathThorn. Although the directory is initialized when the thorn is cloned using GitComponents, the files do not exist. When the build script decides that FUKA must be built from scratch, the script will first pull the FUKA distribution using git. Afterwards, the files are stored in a temporary build location to build the FUKA library.

It is important to note:

The elliptic solver and the initial data codes are not compiled in this procedure

Once the build procedure is finished, only the static library and equations of state are retained

3 Using This Thorn

This thorn is not used other than to ensure the FUKA library is available for the KadathImporter thorn

If one utilises the equations of state stored in HOMExKADATH/eos, the HOMExKADATHxenvironment variable must be correctly defined in the environment of the computer processing the initial data.

3.1 Obtaining This Thorn This thorn can be obtained from [3]

3.2 Support and Feedback The current active developer and maintainer is

xxxxSamuelxTootlex-xtootle@itp.uni-frankfurt.de

4 History

The pre-ETK inclusion KadathThorn and KadathImporter thorns were developed by L. Jens Papenfort, Samuel Tootle, and Elias Most for binary black hole and binary neutron star initial data. Samuel Tootle later rewrote the thorns to encapsulate all initial data available from FUKA, updated to meet ETK inclusion standards, and wrote the documentation.

4.1 Acknowledgements

The authors are grateful to the continued support by Philippe Grandclement, author of the KADATH spectral library, throughout the development of the FUKA codes.

References

[1]xxx”Papenfort, L. Jens et al”, New public code for initial data of unequal-mass, spinning compact-object binaries, ”10.1103/PhysRevD.104.024057”

[2]xxxFrankfurt University/KADATH (codes), ”https://bitbucket.org/fukaws/fuka”

[3]xxxFrankfurt University/KADATH (workspace), ”https://bitbucket.org/fukaws/”

5 Parameters

6 Interfaces

General Implements: kadath

7 Schedule This section lists all the variables which are assigned storage by thorn Fuka/KadathThorn. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE